OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. 4x + 13y = 40 4x + 3y = -40

OpenStudy (smilewithmehun):

4x + 13y = 40 4x + 3y = -40 Multiply one set by -1 -1(4x + 3y = -40) -4x -3y = 40 Then do the same as the last problem of adding the two sets 4x + 13y = 40 -4x - 3y = 40 10y = 40 divide by ten y = 4 The use that answer to substitute for y in one set to find the x value like 4x + 13(4) = 40 4x + 52 = 40 4x = -12 x = -3

OpenStudy (smilewithmehun):

*Then

OpenStudy (anonymous):

thanks, again

OpenStudy (clalgee):

Welcome to Openstudy!